OpenStudy (anonymous):

What is the equation of the line that passes through the points (−2, 3) and (2, 7)? x − y = −1 x − y = −2 x − y = − 5 x − y = − 6

OpenStudy (tygrr321):

gimme a sec

OpenStudy (tygrr321):

m = ( y2 - y1 ) / ( x2 - x1 ) = ( 5 - 3 ) / (-2 - 1 ) = ?

OpenStudy (tygrr321):

@redskychicken u there?

OpenStudy (anonymous):

@tygrr321 yeah sorry I had to afk quick

OpenStudy (tygrr321):

lol ok so u got it ^

OpenStudy (anonymous):

the answer would be negative 0.6 repeating but that is not an answer @ tygrr321

OpenStudy (tygrr321):

its not the answer but it is the start

OpenStudy (anonymous):

oh it's 4

OpenStudy (tygrr321):

wat u mean its 4?

OpenStudy (tygrr321):